# Dinkel: Testing Graph Database Engines via State-Aware Query Generation

### v2

@@ Abstract.
Graph database management systems (GDBMSs) store and manipulate graph data and form a core part of many data-driven applications. To ensure their reliability, several approaches have been proposed to test GDBMSs by generating queries in Cypher, the most popular graph query language. However, Cypher allows queries with complicated state changes and data dependencies, which existing approaches do not support and thus fail to generate valid, complex queries, thereby missing many bugs in GDBMSs.
In this paper, we propose a novel state-aware testing approach to generate complex Cypher queries for GDBMSs. Our approach models two kinds of graph state, query context and graph schema. Query context describes the available Cypher variables and their corresponding scopes, whereas graph schema summarizes the manipulated graph labels and properties. While generating Cypher queries, we modify the graph states on the fly to ensure each clause within the query can reference the correct state information. In this way, our approach can generate Cypher queries with multiple state changes and complicated data dependencies while retaining high query validity. We implemented this approach as a fully automatic GDBMS testing framework, Dinkel, and evaluated it on three popular open-source GDBMSs, namely Neo4j, RedisGraph, and Apache AGE. In total, Dinkel found 60 bugs, among which 58 were confirmed and 51 fixed. Our evaluation results show that Dinkel can effectively generate complex queries with high validity (93.43%). Compared to existing approaches, Dinkel can cover over 60% more code and find more bugs within the 48-hour testing campaign. We expect Dinkel’s powerful test-case generation to benefit GDBMS testing and help strengthen the reliability of GDBMSs.
@@ 1. Introduction
Graph database management systems (GDBMSs) are crucial for modern interconnected, data-driven computer software. By storing data in graph structures, GDBMSs efficiently manage data through graph properties (e.g., pairs of data connected through certain relationships). Because of their high efficiency, GDBMSs have been widely adopted in large language models (Neo4j and Generative AI, 2023; Unifying LLMs and Knowledge Graphs for GenAI: Use Cases and Best Practices, 2023), recommendation systems (White Paper: Powering Real-Time Recommendations with Graph Database Technology., 2024), social networking (NBC News Analyzes Hundreds of Thousands of Russian Troll Tweets Using Neo4j., 2021), and other data-driven applications (Creating a Modern Solution to Finding What’s Lost with Neo4j., 2024; Novartis Captures the Latest Biological Knowledge for Drug Discovery., 2021). Their practicality has led 75% of the Fortune 100 and all of North America’s top 20 banks to adopt the currently most popular GDBMS, Neo4j (DB-Engines Ranking., 2024; Who Uses Neo4j., 2024).
GDBMSs are rapidly evolving and complex systems (e.g., Neo4j has 1.2M LOC), where bugs are inevitable during their development and maintenance. These bugs are critical as they can corrupt the GDBMSs and lead them to malfunction. The application of error-prone GDBMSs in critical domains may lead to serious consequences if a malicious attack were to abuse a previously unknown GDBMS bug. For example, an attacker can leverage bugs in GDBMSs to disclose confidential data (CVE-2018-1000820., 2018) or run a Denial of Service (DoS) attack by repeatedly crashing the GDBMS server (CVE-2020-35668., 2020).
@@
To improve GDBMS reliability, several testing approaches (Hua et al., 2023; Kamm et al., 2023; Jiang et al., 2023b; Zhuang et al., 2023) have been proposed to find bugs by generating queries in Cypher, the most widely adopted graph query language (openCypher., 2024). As GDBMSs are large-scale, complex systems, effectively testing GDBMSs is difficult and hindered by two key challenges, query generation and test-oracle construction. Specifically, to cover a broader spectrum of functionalities and deeper query-processing logic in GDBMSs, testing approaches need to generate complex queries that involve various Cypher language features and complicated data dependencies. To accurately identify bugs (e.g., logic bugs) triggered by the generated queries, testing approaches need suitable test oracles to validate the correctness of query execution. Query generation and test-oracle construction are orthogonal and equally important for effective GDBMS testing.
@@
However, a significant gap exists in the query generation of existing GDBMS testing approaches. While all approaches focus on integrating test oracles to identify bugs triggered by the generated queries, none systematically models the Cypher query language to improve query generation. As a result, these approaches are limited in generating valid complex queries, i.e., queries involving various clause combinations and complicated data dependencies. Generating such complex queries is challenging because the clauses invoked by Cypher queries can change the graph states of the manipulated graph database. Such state changes are visible in the subsequent clauses. Moreover, the data and variables used by queries have different scopes, depending on the clauses they are involved in. Lacking systematic modeling of these Cypher language features, existing approaches cannot effectively generate complex queries, and thus many critical functionalities and deep logic of GDBMSs are not exercised by the testing campaigns of these approaches. Therefore, many deep bugs (e.g., the bug shown in Figure 3) in GDBMSs are missed by existing approaches.
@@
To generate complex queries for testing GDBMSs, we need to systematically model the Cypher query language. Cypher is a declarative language that allows expressive data querying without describing specific control flows to achieve it, which enables GDBMSs to flexibly select efficient strategies to execute queries. However, different from typical declarative query languages (e.g., SQL), Cypher queries can make changes to the graph database state while executing clauses of the queries, whose effects are visible in subsequent clauses (Cypher Query Language Reference., 2022). For example, a graph node created in an outer CREATE clause can be referenced at other clauses (e.g., DELETE and MERGE) later on in the same query. Such characteristics make it challenging to effectively test GDBMSs. On one hand, neglecting these characteristics tends to make the generated queries simpler, resulting in the GDBMSs not having to handle complicated data dependencies. Therefore, the testing campaign cannot reach the deep logic of GDBMSs. On the other hand, failing to correctly handle such characteristics can make the generated Cypher queries invalid, resulting in many queries being discarded by GDBMSs in early stages (e.g, query parsing). For example, the generated queries become invalid when they reference items that are nonexistent or out-of-scope.
@@
In this paper, we propose, to the best of our knowledge, the first systematic approach to effectively generate complex Cypher queries. Orthogonal to the test-oracle construction (e.g., oracle for detecting logic bugs in GDBMSs), our approach fills the gap in the query generation of GDBMS testing research. The approach models graph states and possible state changes caused by the Cypher queries. It abstracts the states into two categories, query context and graph schemas. Query context contains information related to temporary variables declared in the queries (e.g., the type and scope of each variable), while graph schemas maintain the current graph information, including the graph labels and properties. As query context and graph schemas may change at different Cypher clauses, we must update these abstractions while generating Cypher queries. To this end, we propose state-aware query generation. Instead of determining query skeletons for later expressions complementary, our approach incrementally constructs clauses for the generated queries. When constructing a new clause, our approach references only the elements that are accessible with respect to the current query context and graph schema. After the clause is completed, the approach accordingly updates the state information. In this on-the-fly way, our approach can accurately maintain the dynamically evolving graph state and thus effectively generate queries involving complicated data dependencies and state changes.
Based on our approach, we implemented Dinkel, a fully automatic GDBMS testing framework. As Dinkel is designed to improve query generation, it does not integrate advanced test oracles for finding logic bugs, which is orthogonal to our research. Instead, Dinkel employs only ASan alarms, error messages, and code assertions to identify bugs but already outperforms the state-of-the-art. We have evaluated our tool on three popular GDBMSs (i.e., Neo4j (Neo4j. Graphs for Everyone., 2024), RedisGraph (RedisGraph. A graph database as a Redis module., 2023), and Apache AGE (Apache AGE. Graph database optimized for fast analysis and real-time data processing., 2024)). In the evaluation, Dinkel efficiently generated complex Cypher queries and kept a high validity rate (93.43%). Using these queries, Dinkel found 60 unique, previously unknown bugs. So far, 58 of these bugs have been confirmed, and 51 fixed. Many bugs are long-latent and missed by all existing approaches in their extensive evaluation. Compared to existing approaches, Dinkel can cover over 60% more code and find more bugs within the 48-hour testing campaign, owing to the complex Cypher queries generated by Dinkel. These results demonstrate the effectiveness of Dinkel in generating complex and valid Cypher queries and finding bugs in real-world GDBMSs.
In summary, we make the following technical contributions:
  • Novel approach: We abstract the state information of Cypher queries into query context and graph schema. Based on the abstraction, we propose state-aware query generation, which generates Cypher queries by incrementally constructing Cypher clauses and continuously updating the abstracted states.
  • Practical realization: Based on our approach, we realize a fully automatic and practical testing framework, Dinkel, to find bugs in GDBMSs by generating complex Cypher queries.
  • Promising results: We evaluated Dinkel on three popular open-source GDBMSs, namely Neo4j, RedisGraph, and Apache AGE. Dinkel found 60 unique and previously unknown bugs, among which 58 have been confirmed and 51 have been fixed, demonstrating its effectiveness.
2. Background and Motivation
@@
Graph Database Management Systems. Graph database management systems (GDBMSs) utilize graph models to represent the stored data. The most widely used graph model in GDBMSs is the labeled property graph model (Jiang et al., 2023b), which stores interconnected data using nodes connected via relationships (i.e., directed edges between nodes) (openCypher., 2024). Nodes and relationships, so-called graph entities, can be associated with labels and properties. Labels are used to group and classify elements, whereas properties are made up of key-value pairs, providing attribute information of elements.
Figure 1. A property graph modeling a person named Alex owning a car.
Figure 1 shows a simple example represented in a labeled property graph. This example shows a person named Alex owning a vehicle of the type car since 2020. The node on the left-hand side has a label Person and a property Name with a value Alex. This node is connected to another node via a direct edge, i.e. relationship, whose label is OWNS. In this relationship, a property with a key since and a value 2020 is held. The node on the right-hand side has attached the label Vehicle and contains a property named type with value car.
@@
Cypher Query Language. Cypher is the most widely adopted query language for property graph databases (openCypher., 2024) and was proposed by Neo4j, the most popular GDBMS (DB-Engines Ranking., 2024). It is designed as a declarative query language, which allows users to specify the required data without realizing the detailed procedures. The general way to specify data in Cypher is to concretize the graph patterns, which follow the format (n)-[r]->(m) and can be used to reference graph entities that satisfy specified conditions. For example, Figure 2 shows a Cypher query that creates the graph in Figure 1. The query concretizes a graph pattern following the CREATE clause to specify the graph entities to be created.
1CREATE (p:Person {Name: 'Alex'})
2 -[:OWNS {since: 2020}]->
3 (v:Vehicle {type: 'car'});
Figure 2. A Cypher query creating the graph in Figure 1.
@@
Different from another declarative query language, SQL, Cypher does not differentiate between data declaration (DDL), manipulation (DML), and query (DQL) languages. Instead, a Cypher query can create, read, and modify data in a single statement, thereby allowing statements of procedural nature and non-trivial query state manipulation. Generally, Cypher queries access or operate on graphs via clauses, and each query can contain multiple clauses.
@@
Table 1 shows the context-free grammar of Cypher query language. A Cypher query consists of a sequence of clauses. Each clause can be either a reading clause, writing clause, reading/writing clause, projecting clause, or others (e.g., system configuration clause) (Cypher Query Language Reference., 2022). Reading clauses (i.e., MATCH) query the GDBMSs to extract information without modifying the graph entities. Writing clauses (e.g., CREATE, DELETE) modify the data stored in GDBMSs by changing the nodes or relationships in the graph. Reading/writing clauses (e.g., MERGE) can both read and write data in the graph. Projecting clauses (e.g., WITH, RETURN) define expressions to be referenced in the subsequent clauses or the result set. When executing a query with multiple clauses, GDBMSs process these clauses sequentially, and the graph state may change after each clause is processed.
1MERGE (x)<-[:A]-(x)<-[:A]-(x)<-[y:A]->(x)
2DELETE y
3CREATE (x)<-[:B]-()
4DELETE y;
Figure 3. A Cypher query triggering an assertion (”i ¡ n”) failure in RedisGraph.
Figure 3 shows a Cypher query containing multiple clauses. This query triggers an assertion failure in RedisGraph. Figure 4 shows the corresponding graph state changes of Figure 3 after each clause is executed by RedisGraph. The first clause is a MERGE clause that will create graph entities following the specified graph pattern if no graph entity matches the pattern. After this clause is executed, a node x and three relationships with labels A are created. Each relationship is from node x to node x. The subsequent DELETE checks whether the graph entities corresponding to y exist and deletes the existing ones. This clause deletes a relationship. The second CREATE clause creates a new node connected to node x with a new relationship whose label is B. The last clause, DELETE, tries to delete relationship y, but as y has already been deleted and is non-existing, this DELETE clause should do nothing. However, RedisGraph mistakenly recognized y as non-deleted and performed a deletion on a non-existing relationship, which triggered an assertion failure. Such a mistake is caused by the entity ID reuse mechanisms of RedisGraph, which reassigns the ID of the deleted relationship y to the new relationship created by the CREATE clause. When referencing y in the last DELETE clause, RedisGraph mistakenly considers y to still exist because its entity ID is being used.
Cypher queries with multiple clauses changing graph states are commonly used in the real world due to the complex nature of data relationships within graph databases. However, automatically generating such queries (e.g., the query in Figure 3) poses challenges, because (1) the generation needs to be aware of the graph state changes caused by each clause (e.g., MERGE and DELETE); and (2) the intermediate data (e.g., variables x and y) for concretizing these state changes need to be correctly referenced at proper clauses.
@@
Limitation of Existing Approaches. Several approaches have been proposed to test GDBMSs using Cypher (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zhuang et al., 2023; Mang et al., 2024). All of them focus on test oracles to identify bugs triggered by generated queries. They employ only simple template-based query generation, without systematically considering the state changes caused by Cypher clauses. Therefore, their generated queries tend to be structurally simple (i.e., containing fewer clauses) and semantically straightforward (i.e., involving fewer data dependencies). One of the approaches, GDsmith (Hua et al., 2023), integrates techniques to improve the generation of MATCH clauses. It first generates proper graph patterns and WHERE conditions in the MATCH. Then, GDsmith builds a query skeleton according to its grammar template and fills its generated data components (i.e., graph patterns and WHERE conditions). This approach is limited because (1) its query generation is customized for only MATCH clause and cannot be scaled to other clauses; and (2) it still involves no insight for handling the graph state changes caused by Cypher clauses. Therefore, similar to existing approaches, GDsmith is still limited in generating complex queries containing multiple clauses and complicated data dependencies.
Without a systematic model for handling the graph state changes, existing approaches are limited in generating complex Cypher queries. However, many bugs are hidden in the deep logic of DBMS implementations, which may only be covered by complex queries (Jiang et al., 2023a, c). To improve the reliability and security of GDBMSs, proposing an effective approach for generating complex Cypher queries has become imperative and critical for GDBMS testing research.
3. State-Aware GDBMS Testing
In this paper, we propose a novel, state-aware approach for testing GDBMSs using complex Cypher queries. Figure 5 shows the approach overview. Our approach introduces two abstractions, query context and graph schema, to precisely model the graph states maintained by GDBMSs during query processing. Specifically, query context describes the temporary variables declared at each clause along with the corresponding scope and type. Graph schema stores available graph labels and properties at each clause. To generate complex queries that involve multiple clauses, we need to accurately handle the possible state changes caused by these clauses. We propose state-aware query generation, which incrementally constructs Cypher queries clause by clause. Each time a new clause is being constructed, our approach checks the current query context and graph schema, and builds the clause with the available data references. After the clause is constructed, the approach updates the state information accordingly. In this way, our approach can continuously construct Cypher clauses for generating complex queries while accurately maintaining the graph state information for involving complicated data dependencies in the generated queries.

### v3

# Dinkel: State-Aware and Granular Framework for Validating Graph Databases

@@ Abstract.
Graph database management systems (GDBMSs) have been powering many data-driven applications. To ensure GDBMS reliability, several testing approaches have been proposed. However, they all suffer from two key limitations: (1) insufficient support for generating complex and valid queries to exercise deep GDBMS code, and (2) lack of general oracles to validate the execution correctness of arbitrary queries.
In this paper, we propose a novel and practical approach Dinkel, to thoroughly test GDBMSs. Our approach consists of two core techniques. First, to generate complex and valid queries, we model two kinds of graph state, query context and graph schema, to describe the Cypher variables and the manipulated graph labels and properties. We generate queries clause-by-clause, and modify the graph states on the fly to ensure each clause references the correct state information. Second, to generally validate query results, we introduce two fine-grained query transformations: clause-level and expression-level transformations. These transformations can operate on arbitrary queries while preserving their semantics. Dinkel validates GDBMSs by checking whether the transformed query produces the same results as the original. We evaluated Dinkel on three well-known GDBMSs. In total, we found 127 bugs, among which 113 were confirmed, 84 were fixed, and 33 were logic bugs. Compared to existing approaches, Dinkel can cover over 70% more code and find substantially more bugs within a 48-hour testing campaign. We expect Dinkel’s powerful bug detection to lay a practical foundation for GDBMS testing.
@@ 1. Introduction
Graph database management systems (GDBMSs) are crucial for modern interconnected, data-driven computer software (Creating a Modern Solution to Finding What’s Lost with Neo4j., 2024; Novartis Captures the Latest Biological Knowledge for Drug Discovery., 2021; Neo4j and Generative AI, 2023; NBC News Analyzes Hundreds of Thousands of Russian Troll Tweets Using Neo4j., 2021; Yin et al., 2023; Chen and Qian, 2023; Xu et al., 2019). Notably, 75% of the Fortune 100 and all of North America’s top 20 banks have adopted the currently most popular GDBMS, Neo4j (DB-Engines Ranking., 2024; Who Uses Neo4j., 2024). While rapidly evolving and becoming more complex (e.g., Neo4j has 1.2M LOC), GDBMSs inevitably suffer from bugs introduced during their development. These bugs are critical as they can corrupt the GDBMSs and lead to serious consequences. One of the most critical types of bugs are logic bugs, which can silently cause GDBMSs to produce incorrect query results.
@@
To improve GDBMS reliability, testing approaches (Hua et al., 2023; Kamm et al., 2023; Jiang et al., 2023b; Zhuang et al., 2023) have been proposed to find bugs by generating queries in Cypher, the most widely-adopted graph query language (openCypher., 2024). As large-scale, complex systems, GDBMSs are difficult to test effectively due to two key challenges: query generation and test-oracle construction. To cover a broad spectrum of functionalities and deep system logic, approaches need to generate complex queries involving various Cypher language features and complicated data dependencies. To identify bugs, especially logic bugs that cause GDBMSs to silently produce incorrect results for queries, testing approaches require test oracles that can be generally applied to validate arbitrary queries.
However, both of these challenges remain unsolved. All existing approaches focus on proposing new test oracles, but struggle to systematically generate complex and valid queries, leaving much GDBMS logic and code inadequately tested or completely unexercised. Although existing approaches focus on test oracles to identify logic bugs, none of the proposed oracles can generally be applied to validate arbitrary queries. As a result, many logic bugs in GDBMSs get missed, even though they can be triggered by existing queries. The following discusses the details of these two challenges.
@@
Challenge 1: Generating Complex Queries.   In DBMSs, many bugs can only be triggered by complex queries (Jiang et al., 2023a; Jiang and Su, 2024; Jiang et al., 2023c). A complex Cypher query can involve various clause combinations and complicated data dependencies. Generating such complex queries is challenging because the clauses invoked by Cypher queries can change the graph states of the manipulated graph database during their execution. Such state changes are visible in subsequent clauses. Moreover, the data and variables used by Cypher queries have different scopes, depending on the clauses they are involved in. Due to the lack of systematic modeling of these Cypher language features, existing approaches cannot effectively generate complex queries, and thus, many critical functionalities and deep logic of GDBMSs are not exercised by the testing campaigns of these approaches. Therefore, many deep bugs (e.g., the bugs shown in Figure 1 and Figure 2) in GDBMSs are missed by existing approaches.
@@
Modeling Cypher language features is not trivial. Cypher is a declarative language that allows expressive data querying without describing specific control flows, which enables GDBMSs to flexibly select efficient strategies to execute queries. However, different from typical declarative query languages (e.g., SQL), Cypher queries can make changes to the graph database state while executing Cypher clauses of the queries, whose effects are visible in subsequent clauses (Cypher Query Language Reference., 2022). For example, a graph node created in an outer CREATE clause can be referenced by subsequent DELETE clauses. Such characteristics make it challenging to effectively test GDBMSs. On the one hand, neglecting these characteristics tends to make the generated queries simpler, resulting in the GDBMSs not having to handle complicated data dependencies and therefore not allowing the testing campaign to reach the deep logic of GDBMSs. On the other hand, failing to correctly handle such characteristics can make the generated Cypher queries invalid, resulting in many queries being discarded by GDBMSs in the early stages (e.g., query parsing). For example, the generated queries may become invalid when referencing nonexistent or out-of-scope items.
MERGE (x)<-[:A]-(x)<-[:A]-(x)<-[y:A]->(x)
DELETE y
CREATE (x)<-[:B]-()
DELETE y;
Figure 1. Assertion failure found by Dinkel in RedisGraph.
Challenge 2: General Test Oracle.   Existing approaches use query transformations to validate the correctness of query results (Kamm et al., 2023; Jiang et al., 2023b; Zhuang et al., 2023; Mang et al., 2024a). Given a generated Cypher query, they transform it into another query and check whether the results of the transformed query and the original query follow specified relationships. However, these approaches cannot be generally applied to arbitrary queries because they require their generated queries to follow specific query patterns so that their transformations can be applied. For example, GraphGenie (Jiang et al., 2023b) requires the predicate in MATCH clauses to satisfy the preconditions of their transformation rules (e.g., one rule requires the predicates to involve a cyclic graph). Moreover, GraphGenie can transform and validate only queries that contain a single MATCH clause followed by a RETURN. GRev (Mang et al., 2024a) integrates an abstract syntax graph (ASG) for interpreting the predicate in MATCH statements and can thus transform the predicate in more diverse ways and with fewer restrictions. However, the predicate modeling in GRev can only be applied to MATCH statements, while queries involving other common clauses like MERGE, UNWIND, and FOREACH remain unsupported and thus untested.
Such inapplicabilities arise from the coarse-grained transformation used by existing approaches. Specifically, their transformations are at the query level. When transforming a query to another related query, these approaches need to take into account the semantics of the entire query and choose transformation rules according to these parsed semantics. They cannot handle semantics that do not fit their transformation rules (e.g., both GraphGenie and GRev do not support FOREACH clauses). Such coarse-grained transformations make existing approaches inapplicable to various queries that do not match specific query patterns. Therefore, many logic bugs are missed. For example, existing approaches miss the bug shown in Figure 2 because the original query is too complex and does not follow any specified patterns.
// Original: 2 rows
RETURN 0 AS n1
UNION CALL {
FOREACH (n2 IN [] |
FOREACH (n3 IN [] |
MERGE (:A {u:0})))
} MATCH (y)
RETURN y AS n1
UNION CREATE ()
RETURN 0 AS n1
// Transformed: 1 row
FOREACH (n0 IN [] |
FOREACH (x IN [] |
SET x = {}))
RETURN 0 AS n1
UNION CALL {
FOREACH (n2 IN [] |
FOREACH (n3 IN [] |
MERGE (:A {u:0})))
} MATCH (y)
RETURN y AS n1
UNION CREATE ()
RETURN 0 AS n1
Figure 2. Queries exposing a logic bug in Neo4j.
In this paper, we propose a novel and practical approach, Dinkel, which is designed to lay a practical foundation for testing GDBMSs. Dinkel integrates two technical solutions to address the two challenges discussed above.
@@
Solution 1: State-Aware Query Generation.Dinkel models graph states and state changes caused by Cypher queries. It abstracts the states into two categories, query context and graph schemas. Query context contains information related to temporary variables declared in the queries (e.g., the type and scope of each variable), while graph schemas maintain the current graph information, including the graph labels and properties. As query context and graph schemas may change at different Cypher clauses, we must update these abstractions while generating Cypher queries. To this end, we propose state-aware query generation. Instead of determining query skeletons for later expressions complementarily, our approach incrementally constructs clauses for the generated queries. When constructing a new Cypher clause, our approach references only the accessible elements according to the current query context and graph schema. After the clause is completed, the approach updates the corresponding state information. In this on-the-fly way, our approach can accurately maintain the dynamically evolving graph state and thus generate queries involving complicated data dependencies and state changes.
Solution 2: Fine-Grained Query Transformation.   Our idea to validate Cypher queries is inspired by EET (Jiang and Su, 2024), which is designed for SQL queries in relational DBMSs (RDBMSs). EET proposes to validate SQL queries by transforming them at the level of expressions. In this paper, we further generalize this method to fine-grained query transformation, without limiting the transformation units to just expressions. Specifically, instead of analyzing the semantics of a whole query, our approach transforms queries by their fine-grained units. In the context of Cypher queries, the fine-grained units can not only be expressions, but also clauses, which can flexibly change graph states. By operating on these low-level units, we focus on the fine-grained semantics of queries, without the need to analyze the overall query-level semantics. For example, we can transform the first RETURN clause of the original query in Figure 2 by inserting an additional FOREACH clause, which is dead code and has no effects because the array to be iterated over is empty. Therefore, the clauses before and after transformations should be semantically equivalent, and the transformed query should produce the same results as the original one. However, their results differ, indicating that a logic bug was triggered. In this process, Dinkel only needs to ensure the semantic equivalence between clauses, without understanding the semantics of the whole query.
@@
Regarding the transformation for clauses, Dinkel manipulates target clauses to affect the graph state in ways that do not interfere with the semantics of the following clauses (e.g., creating and immediately deleting a node), or introduce dead code (e.g., a MATCH clause matching a node that does not exist) before target clauses. Regarding the transformation for expressions, given an arbitrary expression within clauses, Dinkel can construct a semantically equivalent expression by leveraging logical (e.g., De Morgan’s laws) and algebraic equivalences (e.g., $x\equiv x+0$). Dinkel also uses features specific to the Cypher language for transforming expressions (e.g., x $\equiv$ CASE WHEN true THEN x END).
Ultimately, we implemented Dinkel as a fully automatic GDBMS testing framework. Dinkel will be open-sourced once the paper is ready to be public. We have evaluated our tool on three popular GDBMSs (i.e., Neo4j (Neo4j. Graphs for Everyone., 2024), RedisGraph (RedisGraph. A graph database as a Redis module., 2023), and Memgraph (Memgraph. Open-source graph database, tuned for dynamic analytics environments. Easy to adopt, scale and own., 2024)). In the evaluation, Dinkel efficiently generated complex Cypher queries and kept a high validity rate (89.02%). Using these queries, Dinkel found 127 unique, previously unknown bugs, among which 33 are logic bugs. So far, 113 of these bugs have been confirmed, and 84 fixed. Many bugs are long-latent and missed by all existing approaches during their extensive evaluation. Compared to existing approaches, Dinkel can cover over 70% more code and find more bugs within the 48-hour testing campaign. These results demonstrate that Dinkel significantly outperforms existing approaches in bug detection for GDBMSs.
In summary, we make the following contributions:
  • Novel approach: We tackle two fundamental problems of GDBMS testing via two novel solutions: (1) state-aware query generation, which abstracts the Cypher state information to facilitate query generation; and (2) fine-grained query transformation, which can be generally applied to validate arbitrary Cypher queries by manipulating their clauses and expressions.
  • Practical realization: We realize a fully automatic testing framework, Dinkel, to find bugs in GDBMSs by generating and validating complex Cypher queries.
  • Promising results: We evaluated Dinkel on three popular open-source GDBMSs, namely Neo4j, RedisGraph and Memgraph. Dinkel found 127 unique and new bugs, among which 33 are logic bugs. So far, 113 unique bugs have been confirmed, and 84 fixed. These results demonstrate the effectiveness of Dinkel.
2. Background
@@
Graph Database Management Systems. GDBMSs utilize graph models to represent data. The most widely used graph model is the labeled property graph model (Jiang et al., 2023b), which stores interconnected data using nodes connected via relationships (i.e., directed edges between nodes) (openCypher., 2024). Nodes and relationships, generally referred to as graph entities, can be associated with labels and properties. Labels are used to group and classify elements, whereas properties are made up of key-value pairs, providing attribute information.
@@
Cypher Language. Cypher is the most widely adopted query language for property graph databases (openCypher., 2024) and was proposed by Neo4j. It is designed as a declarative query language, which allows users to specify the required data without realizing the detailed procedures. The general way to specify data in Cypher is to concretize the graph patterns, which follow the format (n)-[r]->(m) and can be used to reference graph entities satisfying specified conditions. For example, the query in Figure 1 concretizes a graph pattern following the CREATE clause to specify the graph entities to be created.
@@
Unlike the query language SQL, Cypher does not differentiate between data declaration (DDL), manipulation (DML), or query (DQL) languages. Instead, a Cypher query can create, read, and modify data in a single query, thereby allowing queries of procedural nature and non-trivial state manipulation. Generally, Cypher queries access or operate on graphs via clauses, and each query can contain multiple clauses.
@@
Table 1 shows the context-free grammar of the Cypher query language. A Cypher query consists of a sequence of clauses. Each clause can be either a reading clause, writing clause, reading/writing clause, projecting clause, or others (e.g., system configuration clause) (Cypher Query Language Reference., 2022). Reading clauses (i.e., MATCH) query the GDBMS to extract information without modifying the graph entities. Writing clauses (e.g., CREATE, DELETE) modify the data stored in GDBMSs by changing the nodes or relationships in the graph. Reading/writing clauses (e.g., MERGE) can both read and write data in the graph. Projecting clauses (e.g., WITH, RETURN) define expressions to be referenced in the subsequent clauses or the result set. When executing a query with multiple clauses, GDBMSs process these clauses sequentially, and the graph state may change after each clause is processed.
@@
Illustrative Example. Figure 1 shows a query with multiple clauses. This query triggers an assertion failure in RedisGraph. Figure 3 shows the corresponding graph state changes during execution of each clause. The first clause, MERGE, will create graph entities following the specified graph pattern if no graph entity matches the pattern. Therefore, a node x and three relationships with labels A are created. Each relationship is from node x to node x. The subsequent DELETE checks whether the graph entity y exists and deletes the existing ones, ultimately deleting one relationship. The second CREATE clause creates a new node connected to node x with a new relationship B. The last clause, DELETE, tries to delete relationship y, but as y has already been deleted, this DELETE should do nothing. However, RedisGraph recognized y as non-deleted and performed a deletion on a nonexistent relationship, which triggered an assertion failure. Such a mistake is caused by the entity ID reuse mechanisms of RedisGraph, which reassigns the ID of the deleted relationship y to the new relationship created by the CREATE clause. When referencing y in the last DELETE clause, RedisGraph mistakenly considers y to still exist because its entity ID is being used.
Cypher queries with multiple clauses changing graph states are commonly used in the real world due to the complex nature of data relationships within graph databases. However, it is extremely challenging to automatically generate and validate such queries, because (1) the generation needs to be aware of the graph state changes caused by each clause (e.g., MERGE and DELETE) and properly reference the intermediate data (e.g., variables x and y); and (2) the validation needs to construct general test oracles in the case of the complicated semantics the query can contain.
3. Approach
We propose Dinkel, a GDBMS testing framework integrating two techniques: state-aware query generation and fine-grained query transformation. Figure 4 shows its overview.
3.1. State-Aware Query Generation
Dinkelintroduces two abstractions, query context and graph schema, to precisely model the graph states maintained by GDBMSs. Based on these abstractions, Dinkel incrementally constructs Cypher queries clause by clause. Each time a new clause is being constructed, our approach checks the current query context and graph schema, builds the clause with the available data references, and updates the state information accordingly. In this way, Dinkel continuously constructs clauses for generating complex queries while accurately maintaining the graph state information to introduce complicated data dependencies.